OpenStudy (anonymous):

Find the value of x so that the line passing through (10, 5) and (x, 9) has a slope of -2

jimthompson5910 (jim_thompson5910):

I'll get you started \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] \[\Large -2 = \frac{9-5}{x-10}\] \[\Large -2 = \frac{4}{x-10}\] \[\Large -2(x-10) = 4\] and I'll let you finish

OpenStudy (anonymous):

x=8?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thank you !

jimthompson5910 (jim_thompson5910):

you're welcome